# An independent search for Jovian neutrinos using BOREXINO data

### v1

@@ IV.1 Searches from other planets using modulation data
We now do the same analysis using the monthly modulation data to search for a signal from Venus, Mars and Saturn, in order to serve as a null test of the analysis technique used for Jupiter. We fit each planet data to Eq. 5, by replacing $\mathcal{R}_{\mathrm{jup}}$ with the corresponding contribution from Venus, Mars, and Sun. The prior used for $\mathcal{R}_{\mathrm{sun}}$ is the same as that used for Jupiter. However, for the corresponding prior on each of the planets we extended the lower limits to negative values in order to test for compatibility with zero flux. For calculating the distance, we used DE442s ephemerides. The marginalized best fit posterior intervals for $\mathcal{R}_{\mathrm{ven}}$, $\mathcal{R}_{\mathrm{mar}}$ and $\mathcal{R}_{\mathrm{sat}}$ can be found in Figs. 11,  12, and  13, respectively. The corresponding best-fit values are tabulated in Table 7. For both Venus, Mars, and Saturn we find the best-fit values for the flux contribution to be less than 0 during the 2011-2013 interval. During 2019-2021, the best-fit values for Mars are consistent with zero flux. We find the best-fit values for Venus flux in the 2019-2021 interval given by $\frac{\mathcal{R}_{\mathrm{ven}}}{(1\,\mathrm{AU})^{2}}=0.17\pm 0.09$ (cpd/100t). Therefore, this is consistent with a non-zero flux at $2\sigma$ c.l. similar to that seen for Jupiter. The inferred flux is comparable to that obtained by using the rate time-series data. For Saturn we find $\frac{\mathcal{R}_{\mathrm{sat}}}{(10\,\mathrm{AU})^{2}}=2.98\pm 1.64$ (cpd/100t) (corresponding to $1.8\sigma$).
We now do hypothesis testing using Bayesian model selection. The null hypothesis is that the monthly modulation data contain a contribution from only Sun, while the alternative hypothesis is that there is an additional contribution from the planet (Venus, Mars, or Saturn). Since we are only interested in a positive flux contribution, we ensure that the lower limits on the prior for $\mathcal{R}_{\mathrm{ven}}$, $\mathcal{R}_{\mathrm{jup}}$, and $\mathcal{R}_{\mathrm{sat}}$ are equal to 0. The Bayes factors for 2011-2013 and 2019-2021 intervals can be found from Table 8. We find that the Bayes factors are much less than one for Mars and Venus for both 2019-2021 and 2011-2013. For Saturn, the Bayes factors are close to 1 for both 2019-2021 and 2011-2013. Therefore, Bayesian model comparison does not support an additional contribution from any additional planet.
@@
Venus $\frac{\mathcal{R}_{\mathrm{ven}}}{(1\,\mathrm{AU})^{2}}$ $\mathcal{U}[-50,50]$ $-0.30^{+0.10}_{-0.10}$ $0.17^{+0.09}_{-0.09}$
Mars $\frac{\mathcal{R}_{\mathrm{mar}}}{(2\,\mathrm{AU})^{2}}$ $\mathcal{U}[-13,13]$ $-0.09^{+0.13}_{-0.13}$ $0.02^{+0.04}_{-0.04}$
Saturn $\frac{\mathcal{R}_{\mathrm{sat}}}{(10\,\mathrm{AU})^{2}}$ $\mathcal{U}[-10,10]$ $-1.04^{+1.80}_{-1.81}$ $2.98^{+1.63}_{-1.64}$
@@ IV.2 Goodness of fit tests with monthly modulation data
V Conclusions
@@
In a recent work, A24 analyzed the BOREXINO 7Be solar neutrino data collected from 2011-2021. This data was found to deviate from the expected trend of $\propto\frac{1}{d_{\mathrm{sun}}^{2}}$. A24 posited that a flux contribution from Jupiter could induce a signal similar to that of the 7Be signal with a time variation consistent with a $\frac{1}{d_{\mathrm{jup}}^{2}}$ dependence and also explain the discrepancy in the observed eccentricity. A24 fit the total BOREXINO rate time series data between 2019-2021 using Bayesian regression to a contribution from Jupiter (in addition to Sun and the radioactive background) and found that Jupiter could contribute upto 6% of the total signal with a significance of $\sim 2\sigma$. A24 then did a similar analysis using the monthly modulation data obtained after subtracting the known values of trend of the data from the rate time-series data between 2011-2013 and 2019-2021. They were able to confirm the previous result for a contribution from Jupiter during both these epochs. This contribution from Jupiter was then argued to result from dark matter particles of mass between 0.1-4 GeV trapped by Jupiter.
@@
In this work, we have independently tried to reproduce this result. We have also done additional tests to ascertain the robustness of this signal. For this purpose, we did exactly a similar search in both the rate time series data and modulation data for a signal from Venus, Mars, and Saturn, to see if a similar analysis shows a null result, since apriori no such signal should be expected from any other planet. One slight difference in our analysis of the monthly modulation data is that we averaged over all the data (cf. Eq. 5) instead of an integral (cf. Eq. 3), as done in A24. We also carried out a hypothesis testing using Bayesian model comparison, by calculating the Bayes factor for an additional contribution from a given planet (Jupiter etc.) compared to the null hypothesis that the data only contain a contribution from the Sun (and possibly the radioactive background, if analyzing the total rate time-series data). Finally, we also carried out $\chi^{2}$ goodness of fit tests using all the hypotheses considered. Our results are as follows:
@@
Therefore, we conclude that even though Bayesian inference shows a flux from Jupiter, about 6% of the 7Be flux during Oct. 2019 to Oct. 2021, Bayesian model comparison provides only marginal evidence for this additional contribution from Jupiter. The Bayesian regression technique used in A24 also show evidence for similar spurious contributions upto $2\sigma$ from Saturn and Venus, although once again Bayesian model comparison does not provide any evidence for this contribution.
Nevertheless, we agree with A24 that additional tests can be done by the BOREXINO collaboration by comparing the direction of neutrino events with respect to Jupiter to obtain an additional degree of freedom.
@@ Acknowledgements.
YHP has been supported by a Summer Undergraduate Research Exposure (SURE) Internship at IIT Hyderabad during summer of 2025. We are thankful to Saeed Ansarifard for patiently explaining all the nuances of the analysis done in A24 and promptly answering all our queries.

### v2

@@ IV.1 Searches from other planets using modulation data
We now do the same analysis using the monthly modulation data to search for a signal from Venus, Mars, and Saturn, in order to serve as a null test of the analysis technique used for Jupiter. We fit each planet data to Eq. 5, by replacing $\mathcal{R}_{\mathrm{jup}}$ with the corresponding contribution from Venus, Mars, and Sun. The prior used for $\mathcal{R}_{\mathrm{sun}}$ is the same as that used for Jupiter. However, for the corresponding prior on each of the planets we extended the lower limits to negative values in order to test for compatibility with zero flux. For calculating the distance, we used DE442s ephemerides. The marginalized best fit posterior intervals for $\mathcal{R}_{\mathrm{ven}}$, $\mathcal{R}_{\mathrm{mar}}$, and $\mathcal{R}_{\mathrm{sat}}$ can be found in Figs. 11,  12, and  13, respectively. The corresponding best-fit values are tabulated in Table 7. For Venus, Mars, and Saturn, we find the best-fit values for the flux contribution to be less than 0 during the 2011-2013 interval. During 2019-2021, the best-fit values for Mars are consistent with zero flux. We find that the best-fit values for Venus flux in the 2019-2021 interval are given by $\frac{\mathcal{R}_{\mathrm{ven}}}{(1\,\mathrm{AU})^{2}}=0.17\pm 0.09$ (cpd/100t). Therefore, this is consistent with a non-zero flux at $2\sigma$ significance, similar to that seen for Jupiter. The inferred flux is comparable to that obtained by using the rate time-series data. For Saturn we find $\frac{\mathcal{R}_{\mathrm{sat}}}{(10\,\mathrm{AU})^{2}}=2.98\pm 1.64$ (cpd/100t) (corresponding to a significance of $1.8\sigma$).
We now do hypothesis testing using Bayesian model selection. The null hypothesis corresponds to the ansatz that the monthly modulation data contain a contribution from only Sun, while the alternative hypothesis is that there is an additional contribution from the planet (Venus, Mars, or Saturn). Since we are only interested in a positive flux contribution, we ensure that the lower limits on the prior for $\mathcal{R}_{\mathrm{ven}}$, $\mathcal{R}_{\mathrm{jup}}$, and $\mathcal{R}_{\mathrm{sat}}$ are equal to 0. The Bayes factors for 2011-2013 and 2019-2021 intervals can be found from Table 8. We find that the Bayes factors are much less than one for Mars and Venus for both 2019-2021 and 2011-2013. For Saturn, the Bayes factors are close to 1 for both 2019-2021 and 2011-2013. Therefore, Bayesian model comparison does not support an additional contribution from any additional planet.
@@
Venus $\frac{\mathcal{R}_{\mathrm{ven}}}{(1\,\mathrm{AU})^{2}}$ $\mathcal{U}[-50,50]$ $0.30^{+0.10}_{-0.10}$ $0.17^{+0.09}_{-0.09}$
Mars $\frac{\mathcal{R}_{\mathrm{mar}}}{(2\,\mathrm{AU})^{2}}$ $\mathcal{U}[-13,13]$ $09^{+0.13}_{-0.13} 0.0$$02^{+0.04}_{-0.04}$
Saturn $\frac{\mathcal{R}_{\mathrm{sat}}}{(10\,\mathrm{AU})^{2}}$ $\mathcal{U}[-10,10]$ $1.04^{+1.80}_{-1.81}$ $2.98^{+1.63}_{-1.64}$
@@ IV.2 Goodness of fit tests with monthly modulation data
V Results after using the full data
V.1 Analysis with rate time-series data
We now do the same analysis as in Sect. III for Jupiter with the event rate time-series data over the full 10 year interval from December 2011 to October 2021. We carry out Bayesian regression on Eq. 1 using the same likelihood and priors as in Sect. III. The marginalized 68% and 90% credible intervals for normal and uniform priors on $\mathcal{R}_{\mathrm{sun}}$ using the full data can be found in Fig. 15 and Fig. 16, respectively for the DE422s ephemerides. These best-fit values for $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ are given by $0.70^{+0.38}_{-0.40}$ and $0.71^{+0.43}_{-0.38}$ (cpd/100t) respectively with significances of $1.75\sigma$ and $1.65\sigma$. The observed flux is about a factor of 2.2 smaller than that obtained using the analysis of 2019-2021 data. Furthermore, the significances are roughly comparable compared to using the data from 2019-2021. We also re-calculated the Bayes factors for the presence of an additional contribution of Jupiter compared to no contribution from Jupiter using the full data. We find Bayes factors of 1.1 and 1.4 for normal and uniform priors on $\mathcal{R}_{\mathrm{sun}}$, respectively. Therefore, for both priors used, the Bayes factors are smaller than those used with the data between 2019-2021 and do not support any additional contribution from Jupiter. We thereby conclude that the results from Bayesian model comparison do not support an additional contribution from Jupiter using the full rate time-series data.
Figure 15: Marginalized 68% and 95% credible intervals for $\mathcal{R}_{\mathrm{B}}$ and $\mathcal{R}_{\mathrm{jup}}$ after normal prior on $\mathcal{R}_{\mathrm{sun}}\in\mathcal{N}(25,2)$ with units of (cpd/100t) using the full 10 years of BOREXINO data. The ephemerides used and the priors on other parameters are same as in Fig. 1. The marginalized 68% value for $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ is given by $0.70^{+0.38}_{-0.40}$ (cpd/100t).
Figure 16: Marginalized 68% and 95% credible intervals for $\mathcal{R}_{\mathrm{sun}}$, $\mathcal{R}_{\mathrm{jup}}$ and $\mathcal{R}_{\mathrm{B}}$ using the full 10 years of BOREXINO data after using uniform prior on $\mathcal{R}_{\mathrm{sun}}\in\mathcal{U}[0,50]$ with units of (cpd/100t). The ephemerides used and the priors on other parameters are the same as in Fig. 1. The marginalized 68% credible interval for $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ is given by $0.71^{+0.43}_{-0.38}$ (cpd/100t).
V.2 Analysis with monthly modulation data
We now do the same analysis as in Sect. IV using the full 10 years of data. We do a Bayesian regression analysis on Eq. 5 using the same priors as in Sect.  IV. The best-fit marginalized values for $\mathcal{R}_{\mathrm{jup}}$ for the two time intervals can be found in Fig. 17 for the DE422s ephemerides. The best-fit value for $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ for the full 10 years of data is equal to $0.65^{+0.42}_{-0.36}$ (cpd/100t) corresponding to a significance of $1.5\sigma$. The observed flux is about 3 times smaller than when using the data from 2019-21. Furthermore, the significance is very slightly reduced compared to using the data from 2011-2013 or 2019-2021. We now calculate the Bayes factor (for an additional contribution from Jupiter) using the full modulation data with the same priors as in Sect.  IV. We find the Bayes factor to be equal to 1.3. Therefore, the Bayes factor is reduced compared to that estimated in the intervals 2011-2013 and 2019-2021 and is consistent with no contribution from Jupiter.
Figure 17: The marginalized posteriors on $\mathcal{R}_{\mathrm{jup}}$ for Jupiter using the monthly modulation binned data (cf. Eq. 5) for the entire 10 years duration. The priors and ephemerides used are same as in Fig. 10. The best-fit value for $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ for the full 10 years of data is equal to $0.65^{+0.42}_{-0.36}$ (cpd/100t).
VI Frequentist analyses of the data
VI.1 Analysis with rate time-series
We now redo this search for a Jovian signal using the rate time-series data by carrying out a frequentist analysis to complement the Bayesian analysis done in the remainder of the manuscript. For this purpose, we use profile likelihood to deal with the nuisance parameters. More details about principle and application of profile likelihood to account to obtain frequentist confidence intervals can be found in  Particle Data Group et al. (2020); Herold et al. (2025) or some of our recent works Ramakrishnan and Desai (2025); Barua et al. (2025). For analyzing the rate time-series data, we kept the solar contribution fixed to $\mathcal{R}_{\mathrm{sun}}/(1\mathrm{AU})^{2}=25$ (cpd/100t). We then construct a grid of values for $\mathcal{R}_{\mathrm{jup}}$. For each value of $\mathcal{R}_{\mathrm{jup}}$, we maximize the combined likelihood $\mathcal{L}(\mathcal{R}_{\mathrm{jup}},\mathcal{R}_{\mathrm{B}})$ with respect to $\mathcal{R}_{\mathrm{B}}$
$\mathcal{L}(\mathcal{R}_{\mathrm{jup}})=\max_{\mathcal{R}_{\mathrm{B}}}\mathcal{L}(\mathcal{R}_{\mathrm{jup}},\mathcal{R}_{\mathrm{B}})$ (6)
[…]
Figure 18: $\Delta\chi^{2}$ as a function of $\frac{R_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ using rate time series data for Oct. 2019 - Oct. 2021. The red dot corresponds to $\Delta\chi^{2}=0$ corresponding to the minimum value of $\chi^{2}$. The dashed horizontal line corresponds to $\Delta\chi^{2}=1$ and the corresponding X-intercept is used to obtain the 68% ($1\sigma$) confidence level estimates of $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$. The best-fit value of $\mathcal{R}_{\mathrm{jup}}$ is given by $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}=1.48\pm 0.8$ (cpd/100t).
The corresponding plot using the full 10 years of BOREXINO data can be found in Fig. 19. Once again, $\Delta\chi^{2}$ shows a parabolic trend as a function of $\mathcal{R}_{\mathrm{jup}}$ with a clear minimum. The best-fit value for $\mathcal{R}_{\mathrm{jup}}$ is given by $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}=0.64\pm 0.43$ (cpd/100t), corresponding to a statistical significance of $1.6\sigma$. This value and significance agree with the Bayesian estimate (cf. Sect. V). Note however that we do not get a central non-zero estimate at 90% c.l., since the $\Delta\chi^{2}=2.71$ value for the full dataset to the left of minimum occurs for $\mathcal{R}_{\mathrm{jup}}<0$. Therefore, we set a 90% upper limit on $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}=<1.3$ (cpd/100t) using the full 10 years of data.
Figure 19: $\Delta\chi^{2}$ as a function of $\frac{R_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ using rate time series data for full 10 years of Borexino data. The plot is structured in the same way as Fig. 18. The best-fit value of $\mathcal{R}_{\mathrm{jup}}$ is given by $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}=0.64\pm 0.43$ (cpd/100t).
Figure 20: $\Delta\chi^{2}$ as a function of $\frac{R_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ using the monthly modulation data for Dec. 2011 - Dec. 2013 (top panel), Oct. 2019 - Oct. 2021 (middle panel), and full 10 years of Borexino data (bottom panel). The plots in all the three panels are structured in the same way as Fig. 18. The central estimate for $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ is equal to $3.12\pm 1.74$ (cpd/100t) and $0.64\pm 0.43$ (cpd/100t), for 2019-2021 and for the full data, respectively, where the uncertainties refer to 68% c.l. These central estimates have been obtained in the same way as in Fig. 18. For 2011-13, we can only set an upper limit of $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}<3.19$ (cpd/100t) at 68% c.l.
VI.2 Analysis with monthly modulation data
We now carry out a similar profile likelihood based frequentist analysis using the monthly modulation data. Since the background is already subtracted from the monthly modulation data, the only nuisance parameter involved is $\mathcal{R}_{\mathrm{sun}}$. Therefore, while analyzing the monthly modulation data, we consider of grid of values for $\mathcal{R}_{\mathrm{jup}}$ and maximize the likelihood (or minimize the $\chi^{2}$) with respect to $\mathcal{R}_{\mathrm{sun}}$ for each value of $\mathcal{R}_{\mathrm{jup}}$. We then find the minimum $\chi^{2}$ value over this grid, and obtain a curve for $\chi^{2}$ as a function of $\mathcal{R}_{\mathrm{jup}}$ similar to the rate time-series data. We do this analysis for three distinct periods: Oct. 2019-Oct. 2021, Dec. 2011-Dec. 2013, and using the full 10 years of data. This plot for all the three periods can be found in Fig. 20. For all the three time intervals, we see parabolic trend for $\Delta\chi^{2}$ as a function of $\mathcal{R}_{\mathrm{jup}}$. However, for Dec. 2011-Dec. 2013 interval, the $\Delta\chi^{2}=1$ value corresponding occurs at $\mathcal{R}_{\mathrm{jup}}<0$. Therefore, we can only set 68% c.l. upper limits for $\mathcal{R}_{\mathrm{jup}}$ for the Dec. 2011-Dec. 2013 given by $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}<3.19$ (cpd/100t). For 2019-2021 and the full 10 years BOREXINO data, we find that $\frac{\mathcal{R}_{\mathrm{jup}}}{(5\,\mathrm{AU})^{2}}$ is equal to $3.12\pm 1.74$ (cpd/100t) and $0.64\pm 0.43$ (cpd/100t), respectively where the uncertainties correspond to 68% c.l. These correspond to significances of $1.8\sigma$ and $1.5\sigma$, respectively. Therefore, except for Dec. 2011- Dec. 2013, these values for $\mathcal{R}_{\mathrm{jup}}$ agree with the Bayesian credible intervals.
VII Conclusions
@@
In a recent work, A24 analyzed the BOREXINO 7Be solar neutrino data collected from 2011-2021. This data was found to deviate from the expected trend of $\propto\frac{1}{d_{\mathrm{sun}}^{2}}$. A24 posited that a flux contribution from Jupiter could induce a signal similar to that of the 7Be signal with a time variation consistent with a $\frac{1}{d_{\mathrm{jup}}^{2}}$ dependence and also explain the discrepancy in the observed eccentricity. A24 fit the total BOREXINO rate time series data between 2019-2021 using Bayesian regression to a contribution from Jupiter (in addition to Sun and the radioactive background) and found that Jupiter could contribute upto 6% of the total signal with a significance of $\sim 2\sigma$. A24 then did a similar analysis using the monthly modulation data obtained after subtracting the known values of trend of the data from the rate time-series data between 2011-2013 and 2019-2021. They were able to confirm the previous result for a contribution from Jupiter during both these epochs. This signal from Jupiter was then argued to result from annihilation of dark matter particles of mass between 0.1-4 GeV trapped in the core of Jupiter.
@@
In this work, we have independently tried to reproduce this result in A24 and also did multiple variants of their analysis. We have also performed additional tests to ascertain the robustness of this signal. For this purpose, we did exactly a similar search in both the rate time series data and modulation data for a signal from Venus, Mars, and Saturn, to see if a similar analysis shows a null result, since apriori no such signal should be expected from any other planet. One slight difference in our analysis of the monthly modulation data is that we averaged over all the data (cf. Eq. 5) instead of an integral (cf. Eq. 3), as done in A24. We also carried out hypothesis testing using Bayesian model comparison, by calculating the Bayes factor for an additional contribution from a given planet (Jupiter etc.) compared to the null hypothesis that the data only contain a contribution from the Sun (and possibly the radioactive background, if analyzing the total rate time-series data). Finally, we also carried out $\chi^{2}$ goodness of fit tests using all the hypotheses considered. Our results are as follows:
@@
  • When we do a Bayesian analysis using the full 10 years of data, the observed flux is about 2 (rate time-series)- 3 (monthly modulation) times smaller than that obtained using only the data from Oct. 2019 - Oct. 2021. However, the values of Bayes factors get reduced to around 1, implying that there is no evidence for additional contribution from Jupiter.
  • We also did a frequentist regression analysis for both the rate time series and monthly modulation data. The results mostly agree with those from Bayesian analysis, except for the monthly modulation data from Dec. 2011 - Dec. 2013, where we only get upper limits at 68% c.l.
Therefore, we conclude that even though Bayesian inference shows a flux from Jupiter, about 6% of the 7Be flux during Oct. 2019 to Oct. 2021, Bayesian model comparison provides only marginal evidence for this additional contribution from Jupiter. This evidence is negligible when analyzing the full 10 years of BOREXINO data. The Bayesian regression technique used in A24 also shows evidence for similar spurious contributions upto $2\sigma$ from Saturn and Venus, although once again Bayesian model comparison does not provide any evidence for this contribution.
Nevertheless, we agree with A24 that additional tests should be done by the BOREXINO collaboration by comparing the direction of neutrino events with respect to Jupiter to obtain an additional degree of freedom. Furthermore, it is also important to check that the statistics of BOREXINO is sufficient to provide the declared sensitivity by anticipating the results with a dedicated Monte Carlo. These studies should be done by the BOREXINO Collaboration. As an additional test, one can also search for Jovian neutrinos using low energy (MeV energy range) data from the Super-Kamiokande experiment which is taking data since 1996.
@@ Acknowledgements.
YHP has been supported by a Summer Undergraduate Research Exposure (SURE) Internship at IIT Hyderabad during summer of 2025. We are thankful to Saeed Ansarifard for patiently explaining all the nuances of the analysis done in A24 and promptly answering all our queries. We are also grateful to the anonymous referee for several constructive comments on our manuscript.